McCann Co. has identified an investment project with the following cash flows.

Year AWN- Cash Flow $850 1,100 1,310 1,170

a. If the discount rate is 9 percent, what is the present value of these cash flows?

b. What is the present value at 16 percent?

c. What is the present value at 27 percent?
Answer #1

$850.00 $1,100.00 $1,310.00 $1.170.00 $779.82 =PV(9%,A2,0,-B2) $925.85 =PV(9%,A3,0,-B3) $1,011.56 =PV(9%A4,0,-B4) $828.86 =PV